# HWC-Loco: A Hierarchical Whole-Body Control Approach to Robust Humanoid Locomotion

### v2

@@ Abstract
Humanoid robots, capable of assuming human roles in various workplaces, have become essential to the advancement of embodied intelligence. However, as robots with complex physical structures, learning a control model that can operate robustly across diverse environments remains inherently challenging, particularly under the discrepancies between training and deployment environments. In this study, we propose HWC-Loco, a robust whole-body control algorithm tailored for humanoid locomotion tasks. By reformulating policy learning as a robust optimization problem, HWC-Loco explicitly learns to recover from safety-critical scenarios. While prioritizing safety guarantees, overly conservative behavior can compromise the robot’s ability to complete the given tasks. To tackle this challenge, HWC-Loco leverages a hierarchical policy for robust control. This policy can dynamically resolve the trade-off between goal-tracking and safety recovery, guided by human behavior norms and dynamic constraints. To evaluate the performance of HWC-Loco, we conduct extensive comparisons against state-of-the-art humanoid control models, demonstrating HWC-Loco’s superior performance across diverse terrains, robot structures, and locomotion tasks under both simulated and real-world environments. Our project page is available at: https://simonlinsx.github.io/HWC_Loco/.
Machine Learning, ICML
@@ 1 Introduction
Humanoid robots, with a physical structure resembling that of humans, can be seamlessly integrated into humans’ workspace and take on roles in completing different tasks (Saeedvand et al., 2019). This capability makes them reliable embodiments of artificial intelligence in various forms (Paolo et al., 2024). In recent years, with advancements in both hardware capabilities and control algorithms, humanoid robots have become an increasingly significant type of robot with a growing impact on practical applications across various environments, such as factories, homes, and offices (Sferrazza et al., 2024; Nasiriany et al., 2024; Tao et al., 2024). However, humanoid robots are typically equipped with complex upper and lower body designs that involve numerous degrees of freedom. Developing robust control algorithms that allow these robots to adapt to a wide range of environments remains a significant challenge.
In the development of humanoid control policy, classic methods commonly rely on model-based optimization (Sakagami et al., 2002; Sentis & Khatib, 2006; Gouaillier et al., 2009; Radford et al., 2015; Chignoli et al., 2021). However, these methods require precise and comprehensive modeling of the robot’s structure, kinematics, and dynamics across different environments. In practice, obtaining such data requires significant computational resources or manual effort. These limitations significantly influence the scalability of these approaches. To develop an end-to-end solution with promising generalizability, recent studies (Zhuang et al., 2024; Radosavovic et al., 2024; Long et al., 2024; He et al., 2024b; Gu et al., 2024b) adopted Reinforcement Learning (RL) methods by training a neural model to control based on human demonstrations and interactions with the environment.
While these learning-based approaches show potential for effective control policies across various tasks, they are typically trained in simulated environments, which often differ considerably from the deploying environment. To bridge the Simulation-to-Real (Sim2Real) gap, previous works often introduce additional regularization to constrain the robot’s movements or employ domain randomization to account for variations in the robot’s physical properties (Lai et al., 2023; Rudin et al., 2022; Nahrendra et al., 2023) . However, excessive regularization can greatly affect the efficiency of control policy, and unstructured randomization often fails to capture safety-critical patterns in real-world applications.
@@
In this paper, to learn a control policy that can dynamically solve the trade-off between maximizing performance and maintaining safety under different deployment environments, we present a Hierarchical Whole-body Control for Robust Humanoid Locomotion (HWC-Loco) that seamlessly incorporates the following policies:
The goal-tracking policy is designed to follow task-specific commands (e.g., moving velocity and direction). Additionally, we ensure that the learned policy can effectively adhere to humanoid behavioral patterns by ensuring its distributional alignment with human motions. Note that this policy primarily focuses on ensuring task completion performance, without addressing safety issues that may arise from potential mismatches with the deployment environment.
The safety-recovery policy aims at ensuring safety across various deployment environments. To address safety-critical scenarios during deployment, we construct an uncertainty set by designing an extreme-case estimation mechanism. Within these constructed scenarios, we propose a dynamic constraint guided by Zero Moment Points (ZMPs) to effectively identify safe control policies. By evaluating whether the ZMPs remain within the support surface, HWC-Loco can assess and ensure the stability of bipedal robots.
While the goal-tracking and safety-recovery policies are primarily designed to enhance task performance and safety control, respectively, they lack a coordination mechanism to effectively control the trade-off between these objectives. To address this limitation, we propose a high-level planning policy that dynamically selects which policy to activate based on the deployment scenario.
We conduct extensive experiments to evaluate the performance of HWC-Loco from four fundamental perspectives: 1) scalability, demonstrated across diverse terrains and robot embodiments, 2) robustness, evaluated under varying scales of environmental disturbances, 3) naturalness, exhibited in its ability to imitate human behavioral norms and movements, and 4) sensitivity reflected in its dynamic adjustment of the low-level policies’ activation rates based on environmental conditions. These validations are performed in both simulated and real-world settings, demonstrating HWC-Loco as a foundational advancement for achieving reliable humanoid locomotion in safety-critical scenarios.
@@ 2 Related Work
Learning-based Legged Robot Locomotion. Various learning-based methods have been proposed for legged locomotion, such as quadrupedal locomotion (Rudin et al., 2022; Shi et al., 2022; Chen et al., 2023; Nahrendra et al., 2023; Margolis & Agrawal, 2023; Zhuang et al., 2023) and bipedal locomotion (Li et al., 2021; Kumar et al., 2022; Li et al., 2023; Duan et al., 2024). Besides, recent studies (He et al., 2024d; Tan et al., 2023) have investigated the design of hierarchical controllers, where a high-level policy guides the lower-level action policy of quadrupedal robots. Despite their success, these approaches cannot be directly applied to humanoid robots because of their complex physical structure and the higher degrees of freedom. Recent advances in hardware and learning methods (e.g., RL algorithms) have enabled humanoid robots to move through diverse and complex environments. Some humanoid control policy navigates across diverse terrains using only proprioceptive information (Radosavovic et al., 2024; Gu et al., 2024c, b). Others enhance performance by incorporating additional sensors, such as vision or LiDAR, to collect detailed environmental data, enabling robots to perform complex tasks like stair climbing and parkour jumps (Zhuang et al., 2024; Long et al., 2024). While these approaches have demonstrated promising locomotion performance, they lack mechanisms to handle safety-critical scenarios that humanoid robots may encounter during real-world deployment.
Whole-body control for humanoid robots. Learning to perform whole-body control remains a critical goal for humanoid robots. By leveraging prior motion trajectories, recent studies (Cheng et al., 2024; He et al., 2024b, a; Fu et al., 2024; He et al., 2024c) have enabled robots to learn human-like movements, such as boxing and dancing. Additionally, previous methods employed multi-stage reward designs to accomplish complex tasks (Kim et al., 2024; Zhang et al., 2024a). These methods have demonstrated the importance of coordinating interactions between the upper and lower body. However, the success of these methods often relies on intricate reward shaping and struggles to generalize across different humanoid robots, particularly when motion priors are unavailable.
@@ 3 Problem Formulation
Partially Observable Markov Decision Process. We formulate the environment of humanoid locomotion task with a Partially Observable Markov Decision Process (POMDP) ${M}=(\mathcal{S},\mathcal{A},\mathcal{O},P_{\mathcal{T}},r,\mu_{0},\gamma)$, where: 1) Within the state space $\mathcal{S}$, a state $s\in\mathcal{S}$ records the complete environmental information and the robot’s internal states. 2) $\mathcal{A}$ denotes the action space, and action $a\in\mathcal{A}$ denotes the target joint angles that a PD controller uses to actuate the degrees of freedom (DOF). 3) $\mathcal{O}$ denotes the partial observations $o\in\mathcal{O}$, which provide incomplete information about the state of the agent due to sensory limitations and environmental uncertainty. At a time step $t$, $\boldsymbol{o}_{t}$ includes velocity command $\boldsymbol{v}_{t}^{cmd}\in\mathbb{R}^{3}$ and proprioception $\boldsymbol{o}_{t}^{p}$ that records a humanoid’s internal state. Appendix A.2 introduces the details. By incorporating the temporal observations $\boldsymbol{o}_{t}^{H}=[\boldsymbol{o}_{t},\boldsymbol{o}_{t-1},\cdots,\boldsymbol{o}_{t-H}]$, the control model can summarize a state as $s_{t}=[\boldsymbol{o}_{t}^{H},\mathbb{P}_{t}]$, where $\mathbb{P}_{t}$ is the privileged information which the robot can’t access in realistic deployment, including base velocity $v_{t}$, terrain height, external disturbance and ZMP features (see Section 4.3). 4) $r$ denotes the reward functions, which typically consist of penalty, regularization, and task rewards. These reward signals significantly influence the optimality of the control policy, for which we provide a detailed introduction in the following. 5) $P_{\mathcal{T}}\in\Delta^{\mathcal{S}}_{\mathcal{S}\times\mathcal{A}}$ denotes the transition function as a mapping from state-action pairs to a distribution of future states. 6) $\mu_{0}\in\Delta^{\mathcal{S}}$ denotes the initial state distribution. 7) $\gamma\in(0,1]$ denotes the discounting factor. The planning stops at a terminating state $\tilde{s}$, and the corresponding terminating time is denoted as $T\in(0,\infty)$.
@@
Under this POMDP, our goal of humanoid locomotion is to learn a policy $\pi\in\Delta^{\mathcal{A}}_{\mathcal{S}}$ according to the objective as follows:
@@
$\mathcal{J}(\pi,{M})=\mathbb{E}_{\mu_{0},p_{\mathcal{T}},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$ (1)
@@
In the task of humanoid locomotion, the total rewards can be generally represented as a weighted combination of task rewards $r_{\mathfrak{T}}$, penalty rewards $r_{\mathrm{P}}$, and regularization rewards $r_{\mathfrak{R}}$ (He et al., 2024b, a; Zhuang et al., 2024):
$r(s_{t},a_{t})=\beta_{\mathfrak{T}}r_{\mathfrak{T}}(s_{t},a_{t})+\beta_{\mathfrak{P}}r_{\mathrm{P}}(s_{t},a_{t})+\beta_{\mathfrak{R}}r_{\mathfrak{R}}(s_{t},a_{t})$
where the rewards have distinct semantic meanings:
@@
  • Task rewards $r_{\mathfrak{T}}$ assess how effectively the agents achieve the goals of the current task, such as tracking velocities (Zhuang et al., 2024), managing contacts (Zhang et al., 2024a), and producing expressive motions (Cheng et al., 2024). These task rewards are closely correlated with the optimality of the current control policy, making them an appropriate objective to maximize in this context.
@@
  • Penalty rewards $r_{\mathrm{P}}$ are used to prevent undesirable events. For example, in locomotion tasks, these penalties are applied when the humanoid falls or when the robot violates dynamic constraints, such as torque limits and joint position limits. To mitigate these negative outcomes, recent studies (He et al., 2024b, a) have assigned a large weight $\beta_{\mathfrak{P}}$ to these rewards. However, we find that these empirical weights inherently function as Lagrange multipliers in constrained optimization problems (Gu et al., 2024a). Consequently, a more effective approach would be to model the control problem as a constrained RL problem.
@@
  • Regularization rewards $r_{\mathfrak{R}}$ are used to align the behavior of the humanoid with human preferences regarding motion styles, safe deployment, and other kinematic or dynamic concerns. However, accurately specifying these regularizations remains challenging since they are often task-dependent, context-sensitive, and inherently rooted in human experts’ experience (Chen et al., 2024). As a result, instead of manually specifying regularizations, we align the robot’s motion with human preferences by imitating human behavior from motion datasets (e.g., AMASS dataset (Mahmood et al., 2019)), which contains rich signals about desirable behavior norms.
@@
$\max_{\pi}\mathcal{J}(\pi,{M})=\max_{\pi}\mathbb{E}_{\mu_{0},p_{\mathcal{T}},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\mathfrak{T}}(s_{t},a_{t})\right]$ (2)
${s.t.}\;\mathcal{D}_{f}(\rho_{{M}}^{\pi^{E}}\|\rho_{{M}}^{\pi})\leq\epsilon_{f}\text{ and }\mathbb{E}_{\tau\sim(\mu_{0},p_{\mathcal{T}},\pi)}\left[\phi(\tau)\right]\geq\epsilon_{\phi}$
@@
Note that we follow previous studies (Moos et al., 2022) and primarily focus on the mismatch in transition dynamics during learning and deployment. The uncertainty set of transition function (Viano et al., 2022) is represented as:
@@
where $\alpha$ specifies the scale of mismatch, $P_{\mathcal{T}}^{L}$ denotes the transition function in the learning environment and $\mathfrak{P}\subseteq\Delta^{\mathcal{{S}}}_{\mathcal{{S}}\times\mathcal{{A}}}$ represents the set of all candidate transition functions. While other factors, such as divergence in state-action spaces and initial state distribution $\mu_{0}$, can influence the performance during deployment, we find that the majority of factors affecting Sim2Real performance in humanoid locomotion tasks can be characterized by the mismatch in transition dynamics. For example, if $\bar{P_{\mathcal{T}}}$ denotes a Gaussian function, the underlying sensor noise can be captured by transitions in $\mathfrak{P}_{\alpha}^{L}$. Additionally, if $\bar{P_{\mathcal{T}}}$ denotes a projection from one spatial state to another, the variations of terrains in the deploying environment can be modeled by transitions in $\mathfrak{P}_{\alpha}^{L}$ (Long et al., 2024).
@@
To solve a robust RL problem, previous studies often consider a max-min objective $\max_{\pi}\min_{{M}\in\boldsymbol{\mathcal{{M}}}_{m}}\mathcal{J}(\pi,{M})$ where $\mathcal{J}(\pi,{M})$ denotes a standard RL objective (1). With this objective, while the agent can learn a conservative policy to ensure worst-case control performance, this policy often compromises the control performance, making the humanoid less effective at tracking the given commands. To address this issue, we focus on ensuring a worst-case feasibility constraint, thereby reducing the influence of mismatched POMDP on maximizing task rewards. Accordingly, we update the CRL objective (2) to the robust humanoid locomotion objective represented as follows:
@@
$\max_{\pi}\min_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{P_{\mathcal{T}}}}\in\mathfrak{P}_{\alpha}^{L}}\mathbb{E}_{\mu_{0},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{\mathcal{T}}^{L}},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\mathfrak{T}}(s_{t},a_{t})\right]$ (4)
${s.t.}\;\mathcal{D}_{f}(\rho_{{M}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi^{E}}\|\rho_{{M}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi})\leq\epsilon_{f}\text{ and }\mathbb{E}_{\tau\sim(\mu_{0},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{P_{\mathcal{T}}}},\pi)}\left[\phi(\tau)\right]\geq\epsilon_{\phi}$
@@
where 1) $\mathfrak{P}_{\alpha}^{L}$ represents the set of mismatched dynamics from the learning environment dynamics ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{\mathcal{T}}^{L}}$ (defined in 3) and 2) $\rho_{{M}}^{\pi}(s,a)=(1-\gamma)\pi(a|s)\sum_{t=0}^{\infty}\gamma^{t}p(s_{t}=s|\pi,{M})$ defines the normalized occupancy measure of policy $\pi$ under the environment ${M}$. Intuitively, we aim to learn a robust control policy that ensures feasibility across all mismatched transition functions ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{P_{\mathcal{T}}}}\in\mathfrak{P}_{\alpha}^{L}$, thereby guaranteeing the safety of policy $\pi$. However, for the reward-maximizing and mimic learning objectives, we optimize the policy based on the transition dynamics in the learning environment ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{\mathcal{T}}^{L}}$, rather than concentrating solely on worst-case guarantees.

### v3

@@ Abstract
Humanoid robots, capable of assuming human roles in various workplaces, have become essential to embodied intelligence. However, as robots with complex physical structures, learning a control model that can operate robustly across diverse environments remains inherently challenging, particularly under the discrepancies between training and deployment environments. In this study, we propose HWC-Loco, a robust whole-body control algorithm tailored for humanoid locomotion tasks. By reformulating policy learning as a robust optimization problem, HWC-Loco explicitly learns to recover from safety-critical scenarios. While prioritizing safety guarantees, overly conservative behavior can compromise the robot’s ability to complete the given tasks. To tackle this challenge, HWC-Loco leverages a hierarchical policy for robust control. This policy can dynamically resolve the trade-off between goal-tracking and safety recovery, guided by human behavior norms and dynamic constraints. To evaluate the performance of HWC-Loco, we conduct extensive comparisons against state-of-the-art humanoid control models, demonstrating HWC-Loco’s superior performance across diverse terrains, robot structures, and locomotion tasks under both simulated and real-world environments. Our project page is available at HWC-Loco.
@@ 1 Introduction
Figure 1: An example of recovering from a Hard Kick: The humanoid robot withstands external disturbance by automatically detecting hazardous states and adjusting its motion to regain stability.
Humanoid robots, with a physical structure resembling that of humans, can be seamlessly integrated into humans’ workspace and take on roles in completing different tasks [1]. This capability makes them reliable embodiments of artificial intelligence in various forms [2]. In recent years, with advancements in both hardware capabilities and control algorithms, humanoid robots have become an increasingly significant type of robot with a growing impact on practical applications across various environments, such as factories, homes, and offices [3, 4, 5].
In the development of humanoid control, classic methods commonly rely on model-based optimization [6, 7, 8, 9, 10]. However, these methods require precise and comprehensive modeling of the robot’s structure, kinematics, and dynamics across different environments. In practice, obtaining such data requires significant computational resources or manual effort. These limitations significantly influence the scalability of these approaches. To develop an end-to-end solution with promising generalizability, recent studies [11, 12, 13, 14, 15] adopted Reinforcement Learning (RL) methods by training a neural model to control based on human demonstrations and interactions with the environment.
While these learning-based approaches show potential for effective control policies across various tasks, they are typically trained in simulated environments, which often differ considerably from the deploying environment. To bridge the Simulation-to-Real (Sim2Real) gap, previous works often introduce additional regularization to constrain the robot’s movements or employ domain randomization to account for variations in the robot’s physical properties [16, 17, 18] . However, excessive regularization can greatly affect the efficiency of control policy, and unstructured randomization often fails to capture safety-critical patterns in real-world applications.
@@
In this paper, to learn a control policy that can dynamically solve the trade-off between task performance and maintaining safety under different deployment environments, we present Hierarchical Whole-body Control for Robust Humanoid Locomotion (HWC-Loco). Figure 1 illustrates an example of HWC-Loco, where a high-level planning policy detects hazardous states and seamlessly switches from the goal-tracking policy to a safety-recovery policy.
The goal-tracking policy is designed to follow task-specific commands (e.g., moving velocity). Additionally, we ensure that the learned policy can effectively adhere to humanoid behavioral patterns by ensuring its distributional alignment with human motions. Note that this policy primarily focuses on ensuring task completion performance, without addressing safety issues that may arise from potential mismatches with the deployment environment.
The safety-recovery policy aims at ensuring safety across various deployment environments. To address safety-critical scenarios during deployment, we construct an uncertainty set by designing an extreme-case estimation mechanism. Within these constructed scenarios, we propose a dynamic constraint guided by the Zero Moment Point (ZMP) to ensure stability under various disturbances.
While the goal-tracking and safety-recovery policies are primarily designed to enhance task performance and safety control, respectively, they lack a coordination mechanism to effectively control the trade-off between these objectives. To address this limitation, we propose a high-level planning policy that dynamically selects which policy to activate based on the scenario.
We conduct extensive experiments to evaluate the performance of HWC-Loco from four fundamental perspectives: 1) effectiveness, demonstrated across diverse terrains, 2) robustness, evaluated under varying scales of disturbances, 3) naturalness, exhibited in its ability to imitate human behavioral norms, 4) scalability evaluated in different robot embodiment and motion tracking tasks. These validations are performed in both simulated and real-world settings, demonstrating HWC-Loco as a foundational advancement for achieving reliable humanoid locomotion in safety-critical scenarios.
@@ 2 Related Work
Learning-based Legged Robot Locomotion. Various learning-based methods have been proposed for legged locomotion, such as quadrupedal locomotion [17, 19, 20, 18, 21, 22, 23] and bipedal locomotion [24, 25, 26, 27]. Despite their success, these approaches cannot be directly applied to humanoid robots because of the complex physical structure and the higher degrees of freedom. Recent advances in hardware and learning methods (e.g., RL algorithms) have enabled humanoid robots to move through diverse and complex environments. Some humanoid control policy navigates across diverse terrains using only proprioceptive information [12, 28, 15]. Others enhance performance by incorporating additional sensors, such as vision or LiDAR, to collect detailed environmental data, enabling robots to perform complex tasks like stair climbing and parkour jumps [11, 13, 29]. While these approaches have demonstrated promising locomotion performance, they lack mechanisms to handle safety-critical scenarios that humanoid robots may encounter during real-world deployment.
Whole-body control for humanoid robots. Learning effective whole-body control remains a central challenge for humanoid robotics. Prior works [30, 31, 32] have demonstrated impressive results in simulated avatars, showcasing the feasibility of learning-based control strategies. More recently, studies such as [33, 14, 34, 35, 36, 37, 38, 39] have enabled realistic robots to imitate complex human-like motions, such as boxing and dancing by leveraging motion priors from human demonstrations. Additionally, previous methods employed multi-stage reward designs to accomplish complex tasks [40, 41]. However, the success of these methods often relies on intricate reward shaping and struggles to generalize across different humanoid robots, particularly when motion priors are unavailable.
@@ 3 Problem Formulation
Learning Environment. We formulate the environment of the humanoid locomotion task with a Partially Observable Markov Decision Process (POMDP) $M=(\mathcal{S},\mathcal{A},\mathcal{O},P_{\mathcal{T}},r,\mu_{0},\gamma)$, where: 1) Within the state space $\mathcal{S}$, a state $s\in\mathcal{S}$ records the environmental information and the robot’s internal states. 2) $\mathcal{A}$ denotes the action space, and action $a\in\mathcal{A}$ denotes the target joint angles that a PD controller uses to actuate the degrees of freedom (DOF). 3) $o\in\mathcal{O}$ denotes an observation, which provides partial information about the state of the agent due to sensory limitations and environmental uncertainty. At a time step $t$, $\boldsymbol{o}_{t}$ includes velocity command $\boldsymbol{v}_{t}^{cmd}\in\mathbb{R}^{3}$ and proprioception $\boldsymbol{o}_{t}^{p}$ that records a humanoid’s internal state. Appendix A.2 introduces the details. By incorporating the temporal observations $\boldsymbol{o}_{t}^{H}=[\boldsymbol{o}_{t},\boldsymbol{o}_{t-1},\cdots,\boldsymbol{o}_{t-H}]$, the control model can summarize a state as $s_{t}=[\boldsymbol{o}_{t}^{H},\mathbb{P}_{t}]$, where $\mathbb{P}_{t}$ is the privileged information which the robot can’t access in realistic deployment, including base velocity $v_{t}$, terrain height, external disturbance and ZMP features (see Section 4.3). 4) $P_{\mathcal{T}}\in\Delta^{\mathcal{S}}_{\mathcal{S}\times\mathcal{A}}$ denotes the transition function as a mapping from state-action pairs to a distribution of future states. 5) $r$ denotes the reward functions, which typically consist of penalty, regularization, and task rewards. These reward signals significantly influence the optimality of the control policy, for which we provide a detailed introduction in the following. 6) $\mu_{0}\in\Delta^{\mathcal{S}}$ denotes the initial state distribution. 7) $\gamma\in(0,1]$ denotes the discounting factor. The robot’s policy stops at a terminating state $\tilde{s}$, and the corresponding terminating time is denoted as $T\in(0,\infty)$.
@@
Under this POMDP, our goal is to learn a policy $\pi\in\Delta^{\mathcal{A}}_{\mathcal{S}}$ according to the objective as follows:
@@
$\max_{\pi}\mathcal{J}(\pi,M)=\max_{\pi}\mathbb{E}_{\mu_{0},p_{\mathcal{T}},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\left[\beta_{\mathfrak{T}}r_{\mathfrak{T}}(s_{t},a_{t})+\beta_{\mathfrak{P}}r_{\mathrm{P}}(s_{t},a_{t})+\beta_{\mathfrak{R}}r_{\mathfrak{R}}(s_{t},a_{t})\right]\right]$ (1)
@@
In the task of humanoid locomotion, the total rewards can be generally represented as a weighted combination of task rewards $r_{\mathfrak{T}}$, penalty rewards $r_{\mathrm{P}}$, and regularization rewards $r_{\mathfrak{R}}$ [14, 34, 11], where the rewards have distinct semantic meanings:
@@
  • Task rewards $r_{\mathfrak{T}}$ assess how effectively the agents achieve the goals of the current task, such as tracking velocities [11], managing contacts [41], and producing expressive motions [33]. These task rewards are closely correlated with the optimality of the current control policy, making them an appropriate objective to maximize in this context.
@@
  • Penalty rewards $r_{\mathrm{P}}$ are used to prevent undesirable events. For example, in locomotion tasks, these penalties are applied when the humanoid falls or when the robot violates dynamic constraints, such as torque limits and joint position limits. To mitigate these negative outcomes, recent studies [14, 34] have assigned a large weight $\beta_{\mathfrak{P}}$ to these rewards. While these weights inherently function as Lagrange multipliers [42], they are predefined and not subject to optimization. Consequently, a more effective approach would be to model the control problem as a constrained RL problem.
@@
  • Regularization rewards $r_{\mathfrak{R}}$ aligns the behavior of the humanoid with human preferences on motion styles or safe deployment. However, accurately specifying these regularizations remains challenging since they are often task-dependent, context-sensitive, and inherently rooted in human experts’ experience [43]. As a result, instead of manually specifying regularizations, we align the robot’s motion with human preferences by imitating human behavior from motion datasets (e.g., AMASS dataset [44]), which contains rich signals about desirable behavior norms.
@@
$\max_{\pi}\mathbb{E}_{\mu_{0},p_{\mathcal{T}},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\mathfrak{T}}(s_{t},a_{t})\right]{s.t.}\;\mathcal{D}_{f}(\rho_{M}^{\pi^{E}}\|\rho_{M}^{\pi})\leq\epsilon_{f}\text{ and }\mathbb{E}_{\tau\sim(\mu_{0},p_{\mathcal{T}},\pi)}\left[\phi(\tau)\right]\geq\epsilon_{\phi}$ (2)
@@
Note that we follow previous studies [45] and primarily focus on the mismatch in transition dynamics during learning and deployment. The uncertainty set of transition function [46] is represented as:
@@
where $\alpha$ specifies the scale of mismatch, $P_{\mathcal{T}}^{L}$ denotes the transition function in the learning environment and $\mathfrak{P}\subseteq\Delta^{\mathcal{S}}_{\mathcal{S}\times\mathcal{A}}$ represents the set of all candidate transition functions. While other factors, such as divergence in state-action spaces and initial state distribution $\mu_{0}$, can influence the performance during deployment, we find that the majority of factors affecting Sim2Real performance in humanoid locomotion tasks can be characterized by the mismatch in transition dynamics. For example, if $\bar{P_{\mathcal{T}}}$ denotes a Gaussian function, the underlying sensor noise can be captured by transitions in $\mathfrak{P}_{\alpha}^{L}$. Additionally, if $\bar{P_{\mathcal{T}}}$ denotes a projection from one spatial state to another, the variations of terrains in the deploying environment can be modeled by transitions in $\mathfrak{P}_{\alpha}^{L}$ [13].
@@
To solve a robust RL problem, previous studies often consider a max-min objective $\max_{\pi}\min_{M\in\boldsymbol{\mathcal{M}}_{m}}\mathcal{J}(\pi,M)$ where $\mathcal{J}(\pi,M)$ denotes a standard RL objective (1). With this objective, while the agent can learn a conservative policy to ensure worst-case control performance, this policy often compromises the control performance, making the humanoid less effective at tracking the given commands. To address this issue, we focus on ensuring a worst-case feasibility constraint, thereby reducing the influence of mismatched POMDP on maximizing task rewards. Accordingly, we update the CRL objective (2) to the robust humanoid locomotion objective represented as:
@@
$\max_{\pi}\min_{{\color[rgb]{0,0,1}\widehat{P_{\mathcal{T}}}}\in\mathfrak{P}_{\alpha}^{L}}\mathbb{E}_{\mu_{0},{\color[rgb]{1,0,0}P_{\mathcal{T}}^{L}},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\mathfrak{T}}(s_{t},a_{t})\right]{s.t.}\;\mathcal{D}_{f}(\rho_{M^{{\color[rgb]{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi^{E}}\|\rho_{M^{{\color[rgb]{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi})\leq\epsilon_{f}\text{ and }\mathbb{E}_{\tau\sim(\mu_{0},{\color[rgb]{0,0,1}\widehat{P_{\mathcal{T}}}},\pi)}\left[\phi(\tau)\right]\geq\epsilon_{\phi}$
@@
where 1) $\mathfrak{P}_{\alpha}^{L}$ represents the set of mismatched dynamics from the learning environment dynamics ${\color[rgb]{1,0,0}P_{\mathcal{T}}^{L}}$ (defined in 3) and 2) $\rho_{M}^{\pi}(s,a)=(1-\gamma)\pi(a|s)\sum_{t=0}^{\infty}\gamma^{t}p(s_{t}=s|\pi,M)$ defines the normalized occupancy measure of policy $\pi$ under the environment $M$. Intuitively, we aim to learn a robust control policy that ensures feasibility across all mismatched transition functions ${\color[rgb]{0,0,1}\widehat{P_{\mathcal{T}}}}\in\mathfrak{P}_{\alpha}^{L}$, thereby guaranteeing the safety of policy $\pi$. However, for the reward-maximizing and mimic learning objectives, we optimize the policy based on the transition dynamics in the learning environment ${\color[rgb]{1,0,0}P_{\mathcal{T}}^{L}}$, rather than concentrating solely on worst-case guarantees.